An airplane flies 720 miles with the wind. In the same amount of time, it can fly 630 miles againt the wind. The speed of the plane still air is 225 mph. Find the speed of the wind.
Answer by mananth(16949)

plane speed = 225 mph
wind speed = x mph

speed against wind 225 - x
speed with wind 225 + x

Distance against wind 630 m
Distance with wind 720 m

t=d/r time against = time with
630 / ( 225 - x)= 720/( 225+x )
630 ( 225+ x)= 720 ( 225 -x )
141750 + 630 x = 162000 - 720 x
720 x 630 x = 20250
1350 x = 20250
/ 1350
x = 15 mph wind speed